What is four times a number plus five more than three times the number and it equals 47
Answer by ptaylor(2198)

Let x=the number
Four times the number=4x
Five more than three times the number=3x+5
Sooooo, we are told that when added together they equal 47
4x+3x+5=47 subtract 5 from each side
4x+3x+5-5=47-5 collect like terms
7x=42 divide both sides by 7
x=6-----------------the number
CK
4*6+3*6+5=47
24+23=47
47=47
Hope this helps----ptaylor